if 18 pound of peas are packed in 2/3-pound bags, how many bags will there be altogther.
Found 2 solutions by stanbon, rwm:
Answer by stanbon(75887)

if 18 pound of peas are packed in 2/3-pound bags, how many bags will there be altogether.
-----
# of bags = 18/(2/3) = 9*3 = 27
===================================
Cheers,
Stan H.
===========

Answer by rwm(914)

27
2/3*27=18
2x/3=18
divide by 2
x/3=9
multiply by 3
x=27